OpenStudy (anonymous):

The formula for the amount with simple interest is A = P(1 + rt). Find t if A = 168, P = 150, and r = 0.08. A. 1.5 B. 2 C. 2.4 D. 3

OpenStudy (anonymous):

A. because: A=P(1+rt) 168=150(1+0.08) Distributive Property first 168=150+12t -150-150 -18=12t Divide both sides by 12 and you get 1.5